OpenStudy (anonymous):

Verify that each equation is an identity cscx / (cotx + tanx) = cosx So I rewrote each trig sign as (

OpenStudy (anonymous):

(1/sinx) / (Cosx/sinx + Sinx/Cosx) = cosx So now what?

OpenStudy (anonymous):

add the fractions in the bottom

OpenStudy (anonymous):

I know I should know this, but they don't have the same denominators, so how do I fix that since it is a fraction in a fraction?

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

or in your case, if you call \(\cos(x)=a\) and \(\sin(x)=b\) it is \[\frac{a}{b}+\frac{b}{a}=\frac{a^2+b^2}{ab}\]

OpenStudy (anonymous):

then you have \(a^2+b^2=1\) so the denominator is \[\frac{1}{ab}\] and \[\frac{\frac{1}{b}}{\frac{1}{ab}}=\frac{1}{b}\times \frac{ab}{1}=a\] \

OpenStudy (anonymous):

the only one piece of this that had anything to do with trig is that \[a^2+b^2=\cos^2(x)+\sin^2(x)=1\] the rest is algebra

OpenStudy (anonymous):

Thank you so much, you are a grade saver.